Mrs. Wilson is 14 years younger than her aunt. If Mrs. Wilson's age in years is as much below 60 as her aunts age is over 40, how old is each
Answer by ankor@dixie-net.com(22740)

Mrs. Wilson is 14 years younger than her aunt.
w = a - 14
;
If Mrs. Wilson's age in years is as much below 60 as her aunts age is over 40, how old is each
Let c = amt below 60 and above 40
:
40 + c = a
c = a - 40
and
w = 60 - c
c = 60 - w
Therefore
a - 40 = 60 - w
a = 60 + 40 - w
a = 100 - w
Replace w with (a-14)
a = 100 -(a-14)
a = 100 - a + 14
a + a = 114
a = 114/2
a = 57 yrs is the age of the aunt
then
57 - 14 = 43 yrs is W's age